# Adiabatic theorem for non-Hermitian quantum systems with non-degenerate real eigenvalues: A proof following Kato’s approach

### v1

@@ Abstract
The adiabatic theorem is one of the most interesting and significant theorem in quantum mechanics. In 1950, T. Kato gave an elegant proof of this result [1]. However, the validation of adiabatic theorem for non-Hermitian quantum systems is unrevealed. In this paper, by following Kato’s approach, we prove rigorously that the adiabatic theorem is still valid for non-Hermitian systems with non-degenerate real eigenvalues. Moreover, our proof utilizes the complex Berry phase, instead of the orthogonal projections used in Kato’s work.
@@ I Introduction
An interesting and important issue of time-dependent quantum systems is the adiabatic theorem. The theorem states that if a quantum system evolves slowly enough under a time-dependent non-degenerate Hamiltonian and the initial state is in the instantaneous eigenstate of the Hamiltonian, then the final state of the system will remain in the instantaneous eigenstate up to a multiplicative phase factor [2]. In 1950, T. Kato gave an elegant but technically mathematical proof of the theorem [1]. The theorem has attracted much of the interests, and many related topics were discussed [3, 4]. The adiabatic theorem is also intimately related to the Berry phase, which plays an important role in the fields such as geometric phase and quantum computing [5, 6]. However, it should be mentioned that Kato’s discussions on the adiabatic theorem does not involve the Berry phase [1].
The discussions of non-Hermitian quantum systems has a long history and recently there is a growing interests in such systems, for their novel features and potential applications [7, 8, 9]. In particular, the dynamics and topology of non-Hermitian systems, such as the skin effect, has attracted much attentions [10, 11, 12, 13, 14, 15, 16, 17]. The original discussions of the adiabatic theorem is within the framework of Hermitian systems. Recently there are discussions on the adiabatic theorem and geometric phase in non-Hermitian systems [18, 19]. One conjecture is that the adiabatic theorem may fail for the non-Hermitian systems.
In this paper, we discuss the adiabatic following problem, with the focus on the Hamiltonians with non-degenerate real eigenvalues. By following Kato’s approach, we prove rigorously that the adiabatic theorem is still valid for non-Hermitian systems with non-degenerate real eigenvalues. The structure of the paper is organized as follows. In section II, we briefly review the adiabatic theorem. In section III, it is shown that the adiabatic theorem is always valid for Hamiltonians with non-degenerate real eigenvalues. In section IV, we give some discussions, show the insights for the result, and conclude this paper.
@@ II Brief review of adiabatic theorem
Usually, the adiabatic following problem is formulated in a Hermitian system. We briefly review the discussions of the adiabatic theorem, following the way in [1, 20].
@@
where $|j(s)\rangle$ is the $j$-th eigenstate of $H(s)$ and $P_{j}(s)$ is the corresponding orthogonal projection. Briefly speaking, when the initial state is set in $|j(0)\rangle$ and evolves slowly enough, the theorem or equivalently Eq. (2) guarantees that the state follows the $j$-th eigenstate $|j\rangle$ up to a factor. The proof of the theorem utilizes the properties of orthogonal projections and unitary operators can be found in [1, 20]. Thus when discussing the non-Hermitian quantum systems, the proof is not directly applicable.
@@ III.1 The generalized adiabatic theorem and its proof
The adiabatic theorem is valid for Hamiltonians with non-degenerate real eigenvalues.
Let $H(s)$ be a Hamiltonian with non-degenerate eigenvalues, i.e.
@@
where $\lambda_{i}(s)\neq\lambda_{j}(s)$ if $i\neq j$ and $v_{i}(s)$ the corresponding eigenvectors. Since the eigenvalues are different for any $s$, $H(s)$ is diagonalizable and $v_{i}(s)$ form a basis of the space. Now let $\xi_{i}(s)$ be the corresponding biorthogonal vectors, i.e. $\xi_{i}^{\dagger}(s)v_{i}(s)=\delta_{ij}.$ Note that $\xi_{i}(s)$ always exist. In fact, denote $V(s)$ the matrix whose column vectors are $v_{i}(s)$, then $\xi_{i}^{\dagger}(s)$ are the row vectors of the inverse of $V(s)$. Now $H(s)$ and $I$ can be rewritten as
$H(s)=\sum_{i}\lambda_{i}(s)v_{i}(s)\xi_{i}^{\dagger}(s),$ (4)
$I=\sum_{i}v_{i}(s)\xi^{\dagger}(s).$ (5)
Moreover, for any $j$, one can define a matrix $S(s)$ by
$S(s)=\sum_{i\neq j}\frac{1}{\lambda_{i}(s)-\lambda_{j}(s)}v_{i}(s)\xi_{i}^{\dagger}(s),$ (6)
@@
Moreover,
$\frac{d}{ds}[e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)]$ (11)
$=$ $-e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}\langle\xi_{j}(s)|\dot{v}_{j}(s)\rangle v_{j}(s)$
$+$ $e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}\dot{v}_{j}(s)$
$=$ $e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}\sum_{i\neq j}\langle\xi_{i}(s)|\dot{v}_{j}(s)\rangle v_{i}(s)$
$=$ $[I-v_{j}(s)\xi_{j}^{\dagger}(s)]\frac{d}{ds}[e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)].$
@@
$\frac{d}{ds}[e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)]$
$=$ $[I-v_{j}(s)\xi_{j}^{\dagger}(s)]\frac{d}{ds}[e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)]$
$=$ $(H(s)-\lambda_{j}(s)I)S(s)\frac{d}{ds}[e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)].$
For simplicity, denote $e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}v_{j}(s)$ by $\tilde{v}_{j}(s)$. Eq. (III.1) can be rewritten as
$\frac{d}{ds}[\tilde{v}_{j}(s)]=(H(s)-\lambda_{j}(s)I)S(s)\frac{d}{ds}[\tilde{v}_{j}(s)].$ (13)
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j}(s)]$ (14)
$=$ $\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j}(s)$
$+$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j}(s)]$
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j}(s)$ (15)
$=$ $iTe^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)[H(s)-\lambda_{j}(s)I]\tilde{v}_{j}(s)$
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j}(s)]$ (16)
$=$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j}(s)].$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\tilde{v}_{j}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]d\sigma.$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{s}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)(H(\sigma)-\lambda_{j}(\sigma)I)S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]|_{0}^{s}$
- $\frac{1}{iT}\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j}(\sigma)]]d\sigma.$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j}(s)-\tilde{v}_{j}(0)\rightarrow 0.$ (18)
@@
$U_{T}(s)v_{j}(0)\rightarrow e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle dr}v_{j}(s).$
This is exactly the adiabatic theorem. When the initial state is the eigenstate and the system evolves slowly enough, the final state remains the instantaneous eigenstate up to a factor. The factor comes from two parts, one is the dynamic phase and the other is the Berry phase.

### v2

# The adiabatic theorem for non-Hermitian quantum systems with real eigenvalues and the complex geometric phase

@@ Abstract
The adiabatic theorem is one of the most interesting and significant theorems in quantum mechanics. However, the adiabatic theorem can fail for general non-Hermitian quantum systems. In this paper, by utilizing the complex geometric phase, the functional calculus for biorthogonal systems and the Grönwall inequality, we prove rigorously that the adiabatic theorem is still valid for diagonalizable non-Hermitian systems with real eigenvalues. The proof also justifies the definition of a complex Berry phase in non-Hermitian systems.
@@ I Introduction
An interesting and important issue of time-dependent quantum systems is the adiabatic theorem. The theorem states that if a quantum system evolves slowly enough under a time-dependent non-degenerate Hamiltonian and the initial state is in the instantaneous eigenstate of the Hamiltonian, then the final state of the system will remain in the instantaneous eigenstate up to a multiplicative phase factor [1]. In 1950, T. Kato gave an elegant but technically involved proof of the theorem [2]. Since its discovery, the adiabatic theorem has attracted much interests from researchers, and many related topics were discussed [3, 4, 5]. In particular, it is intimately related to the studies of geometric phase and quantum computing, with a wide range of applications [6, 7, 8]. However, it should be mentioned that the original discussions on the adiabatic theorem does not involve the Berry phase, which is now usually considered as an indispensable companion with the adiabatic evolution. In fact, it is more than fifty years later when Berry systematically investigated the geometric phase bearing his name [6]. Later, the concept of geometric phase is also extended to the non-abelian case [9].
The discussions of non-Hermitian quantum systems has a long history and recently there is a growing interests in such systems, for their novel features and potential applications [10, 11, 12]. In particular, the dynamics and topology of non-Hermitian systems, such as the skin effect, has attracted much attentions [13, 14, 15, 16, 17, 18, 19, 20]. The original discussions of the adiabatic theorem is within the framework of Hermitian systems. Recently there are discussions on the adiabatic theorem and geometric phase in non-Hermitian systems [21, 22, 23, 24, 25, 26, 27, 28]. A key difference between the adiabatic evolution of Hermitian and non-Hermitian systems is that the adiabatic theorem may fail for the latter ones.
In this paper, we discuss the adiabatic following problem, with the focus on diagonalizable Hamiltonians with real eigenvalues. By utilizing a complex geometric phase, the Grönwall inequality and generalizing Kato’s functional calculus techniques to biorthogonal systems, we prove rigorously that the adiabatic theorem is still valid for non-Hermitian systems with real eigenvalues. The structure of the paper is organized as follows. In section II, we briefly review the adiabatic theorem. In section III, it is shown that the adiabatic theorem is always valid for diagonalizable Hamiltonians with real eigenvalues. In section IV, we give some discussions, show the insights for the result, and conclude this paper. Section V is the appendix.
@@ II Brief review of adiabatic theorem
Usually, the adiabatic following problem is formulated in a Hermitian system. We briefly review the discussions of the theorem, following the way in [2, 29].
@@
where $|j(s)\rangle$ is the $j$-th eigenstate of $H(s)$ and $P_{j}(s)$ is the corresponding orthogonal projection. Briefly speaking, when the initial state is set in $|j(0)\rangle$ and evolves slowly enough, the theorem or equivalently Eq. (2) guarantees that the state follows the $j$-th eigenstate $|j\rangle$ up to a factor. On the other hand, in the degenerate case, the state will remain in the subspace spanned by the eigenvectors of the same eigenvalue, which introduced the non-abelian geometric phase [9]. The proof of the adiabatic theorem utilizes the properties of orthogonal projections and unitary operators can be found in [2, 29]. Thus when discussing the non-Hermitian quantum systems, the proof is not directly applicable.
@@ III.1 The generalized adiabatic theorem and its proof
The adiabatic theorem is valid for diagonalizable Hamiltonians with real eigenvalues.
Let $H(s)$ be a Hamiltonian satisfying
@@
where $\lambda_{i}(s)$ and $v_{i}(s)$ are the real eigenvalues and corresponding eigenvectors, respectively. Now let $\xi_{i}(s)$ be the corresponding biorthogonal vectors, i.e. $\xi_{i}^{\dagger}(s)v_{i}(s)=\delta_{ij}.$ Note that $\xi_{i}(s)$ always exist. In fact, denote $V(s)$ the matrix whose column vectors are $v_{i}(s)$, then $\xi_{i}^{\dagger}(s)$ are the row vectors of the inverse of $V(s)$. Since the eigenvalues can be degenerate, $H(s)$ and $I$ can be rewritten as
$H(s)=\sum_{k}\lambda_{k}(s)P_{k}(s),$ (4)
$I=\sum_{k}P_{k}(s),$ (5)
where $\lambda_{k}(s)$ are the different eigenvalues and $P_{k}(s)$ are the projections such that $P_{k}(s)P_{l}(s)=\delta_{kl}P_{k}(s)$. To see the existence of $P_{k}(s)$, note that one can take the eigenvectors of the same eigenvalue together. For example, let the eigenvectors of $\lambda_{k}(s)$ be $v_{k_{j}}(s)$, then $P_{k}(s)=\sum_{k_{j}}v_{k_{j}}(s)\xi^{\dagger}_{k_{j}}(s)$. Moreover, it is obvious that $P_{k}(s)$ is independent of the choice of $v_{k_{j}}(s)$. In fact, if one takes $v^{\prime}_{k_{j}}(s)$ and $\xi^{\prime}_{k_{j}}(s)$, then direct calculations show that $P_{k}(s)=\sum_{k_{j}}v_{k_{j}}(s)\xi^{\dagger}_{k_{j}}(s)=\sum_{k_{j}}v^{\prime}_{k_{j}}(s)(\xi^{\prime})^{\dagger}_{k_{j}}(s)$ since $v_{k_{j}}(s)$, $v^{\prime}_{k_{j}}(s)$ are related by an invertible matrix and $\xi_{k_{j}}(s)$, $\xi^{\prime}_{k_{j}}(s)$ are related by its inverse matrix.
Now for any $j$, one can define a matrix $S(s)$ by
$S(s)=\sum_{k\neq j}\frac{1}{\lambda_{k}(s)-\lambda_{j}(s)}P_{k}(s),$ (6)
@@
On the other hand, there exists some operator $K_{j}(s)$ satisfying
$\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]=[I-P_{j}(s)]\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)].$ (11)
To see this, note that one can define $K_{j}(s)$ by
$\langle\xi_{i}(s)|K^{\prime}_{j}(s)|v_{k}(s)\rangle=-\sum_{l}\langle\xi_{i}(s)|K_{j}(s)|v_{l}(s)\rangle\langle\xi_{l}(s)|\dot{v}_{k}(s)\rangle.$
Equivalently, the above equation can be written in the matrix form
$K^{\prime}_{j}(s)=-K_{j}(s)G(s),$
where $G_{lk}(s)=\langle\xi_{l}(s)|\dot{v}_{k}(s)\rangle$. The above equation has solution which are often represented by using the time ordering operator. Moreover, one can verify that the above operator satisfies Eq. (11). Moreover, $K_{j}(s)v_{j_{l}}(s)$ lies in the subspace of $\lambda_{j}(s)$.
@@
$\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]$
$=$ $[I-P_{j}(s)]\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]$
$=$ $[H(s)-\lambda_{j}(s)I]S(s)\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)].$
For simplicity, denote $K_{j}(s)v_{j_{l}}(s)$ by $\tilde{v}_{j_{l}}(s)$. Eq. (III.1) can be rewritten as
$\frac{d}{ds}[\tilde{v}_{j_{l}}(s)]=(H(s)-\lambda_{l}(s)I)S(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)].$ (14)
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)]$ (15)
$=$ $\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j_{l}}(s)$
$+$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)]$
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j_{l}}(s)$ (16)
$=$ $iTe^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)[H(s)-\lambda_{j}(s)I]\tilde{v}_{j_{l}}(s)$
@@
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)]$ (17)
$=$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)].$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma.$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{s}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)(H(\sigma)-\lambda_{j}(\sigma)I)S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]|_{0}^{s}$
- $\frac{1}{iT}\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]]d\sigma.$
@@
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j_{l}}(0)\rightarrow 0.$ (19)
@@
$U_{T}(s)v_{j_{l}}(0)\rightarrow e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}K_{j}(s)v_{j_{l}}(s).$ (20)
This is exactly the adiabatic theorem. When the initial state is the eigenstate and the system evolves slowly enough, the final state remains in the subspace of instantaneous eigenstates. In particular, when the eigenvalues are non-degenerate, one has
$U_{T}(s)v_{j}(0)\rightarrow e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle dr}v_{j}(s),$
the state remains an eigenstate up to a factor. The factor consists of two parts, one is the dynamic phase and the other is the Berry phase.